OpenStudy (anonymous):

PLEASE HELP WILL FAN AND MEDAL!!! Solve And Check r/6 = 5 2/3 A. 34 B. 30 2/3 C. 20 D. 17/18

OpenStudy (qwertty123):

Hello!

OpenStudy (anonymous):

Hi :) Can You Please Help?

OpenStudy (qwertty123):

Sure

OpenStudy (anonymous):

Thank You. :) So How Do I Solve It?

OpenStudy (qwertty123):

Hold on! I will solve it and then show you!

OpenStudy (anonymous):

First, you should simplify both sides of the equation. Do you know how?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Well, let's simplify the right side of the equation first. In order to simplify the fraction 5 2/3, you multiply 5 by 3 and add 2: 5 x 3 + 2, so you get 17/3. Understand?

OpenStudy (anonymous):

Yes! :)

OpenStudy (anonymous):

Good! So now we have: r/6 = 17/3 Our aim is to get the variable by itself, all alone, on one side. So let's get rid of 6 by multiplying 6 on both sides. Can you do that?

OpenStudy (anonymous):

Do I Multiply 6 x 5?

OpenStudy (anonymous):

No, you just have to multiply 6 on both sides of the equation like this: r/6 (6) = 17/3 (6)

OpenStudy (anonymous):

So 6 x 6?

OpenStudy (anonymous):

The left side should be easy, it is the main reason you are multiplying the equation by 6 since you want to get rid of it so you can get r alone. So the left side becomes r. To solve the right side, you just have to multiply 17/3 by 6: 17/3 x 6

OpenStudy (anonymous):

Do you know how you got rid of 6 on the left side?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

17/3 x 6 Would Be 34 So Would It Be A?

OpenStudy (anonymous):

Yes, that's correct! Your final answer is r = 34. You were able to get rid of 6 on the left side because it was division: r/6. Multiplication cancels out division, like negatives and positives.

OpenStudy (anonymous):

Thank You So Much!

OpenStudy (anonymous):

Glad to help any time! :)